Solve the system for x and y
y -2x = 0
x^2 + y^2 = 5

Answer by ewatrrr(24785)


Hi
solving the system by Substitution
y -2x = 0 Or y = 2x
x^2 + y^2 = 5
x^2 + (2x)^2 = 5
x^2 + 4x^2 = 5
5x^2 = 5
x^2 = 1
x = ± 1
y = ± 2